# Bridging Social Media and Search Engines: Dredge Words and the Detection of Unreliable Domains

### v2

@@ Introduction
Classifying individual texts or articles is a core problem addressed by research in misinformation detection. Such detection methods have typically relied on website content and social media data (Castelo et al. 2019; Chen and Freire 2020; Silva et al. 2021). Castelo et al. proposed a topic-agnostic detection system (Castelo et al. 2019) that identifies unreliable articles based on Linguistic Inquiry and Word Count features. Chen and Freire adopted the method for the task of unreliable domain discovery (Chen and Freire 2020). Their discovery system capitalized on user tendencies within the social graph, wherein a user who tweets a URL from a known unreliable source is likely to also tweet URLs from yet unknown sources. Similarly, Silva et al. combine website content and social context resulting in a misinformation page detection method that leverages heterogeneous data types (Silva et al. 2021), with a focus on early detection across a broad range of topics. The predictive power of webgraphs and social media data have been demonstrated separately in several contexts. Aswani et al. detect SEO manipulation by link-building sites by clustering on Pagerank score and Domain Authority (Aswani et al. 2021). For detecting news site bias, Aires et al. Patricia Aires, G. Nakamura, and F. Nakamura (2019) scrape cross-links from a list of prominent news sites from Media Bias Fact Check (Zandt 2022). Sehgal et al. explore a case where misinformation was spread through coordinated hyperlink and social media networks (Sehgal et al. 2021). Another recent case study demonstrated the manipulated webgraph linkages of unreliable pseudo-thinktanks (Williams and Carley 2023). Additionally, Zhang and Cabage show that social media and SEO promotion have different strengths; they find that social-sharing results in immediate but short-term boosts to traffic, while the benefits of link-building are slower to realize, but last longer (Zhang and Cabage 2017). This hints that a combination of both webgraph and social media data should be used in investigating misinformation sources. However, to our knowledge, this is the first work that attempts to combine the two contexts for unreliable domain classification or discovery.
@@
We use the domain credibility labels published by Lin et al. (2023). The authors of the work align the domain reliability ratings of all domains ranked by 6 expert groups and run imputation followed by principal component analysis to generate aggregate ratings for 11,520 domains. We binarize the news domain rankings as ”reliable” and ”unreliable” using a threshhold of 0.5162, corresponding to the bottom two quintiles of the data. Following the methodology of Carragher, Williams, and Carley (2024a), We use the SEO toolkit service Ahrefs333ahrefs.com to extract the 10 domains which link to each of the 11,520 target domains at the highest volume (the highest-volume back-linking domains). Ahrefs feature values have been found to correlate strongly with ground-truth pagerank calculations and competing traffic estimate systems (Carragher, Williams, and Carley 2024b). In total, we extracted 43,758 domains, each with 23 domain-level attributes. Additional discussion of labels, collection details, and EDA of the webgraph can be found in Appendix A. Each edge in this dataset denotes a direct path, as a user on the source site could click a hyperlink to move to the reliability-labeled target site. We were unable to extract features for 193 domains—the domains largely appeared dead or inactive, so we excluded them from the data. This left us with 11,327 labeled domains.
@@
From this set of 3.6 billion tweets, we extract all Tweets that mention or link to one of our 11,327 websites, resulting in a dataset of 320M tweets that link to over 840k unique domains. Of the original 11,327 labeled domains, only 5,504 appeared within the Twitter data. To reduce the noise in the dataset we undertake several cleaning and filtration operations, which, due to length constraints, we detail in Appendix A. Following these cleaning operations, we are left with 555k users and 4.9M unique tweets which we observed tweeted or retweeted 91.2M times. This reduced dataset contains mentions of 2,475 reliability-labeled domains and 714 unlabeled back-linking domains.
@@
For the 100 websites with the respective lowest and highest reliability rankings, we extract the top-ranked $k\leq$ 1,000 Google keyphrases from Ahrefs, for a total of 34,646 keyphrases. We then used WebSearcher (Robertson and Wilson 2020) to query each of these keyphrases on Google and extracted the first 10 URLs returned in each SERP. We kept all queries for which the target unreliable domain was returned in the top 10 Google search results. As many of the domains for which we collected dredge words did not have any queries for which they ranked in the top 10 on Google, this resulted in a set of 3,939 dredge words spanning 46 unreliable domains. We created a separate query to find mentions of these keyphrases in the 3.6B tweet covid dataset. This yielded 5.7 million tweets containing dredge words. Many of the most common mentions were explicit mentions of domain names or organization names (which can be identical to twitter handles—like “gatewaypundit”, “infowars”, and “nvic” all received over 10,000 mentions. To filter down the data, we use regular expressions to ensure each drudge word begins with a hashtag, starts a tweet, or is preceded by a white space. This retains 213 dredge words in 421k tweets that qualitatively contain less noise. Of these 421k tweets, only 9,788 (2%) explicitly linked to the unreliable domain associated with the dredge word.
@@
The user, who is encountering new information, might then be curious and compelled to search “indigo children meaning” or just “indigo children” on Google—both phrases are dredge words that surface Gaia in the top 10 results. Interestingly, after manually searching these “indigo children meaning” in June 2024, we found that Google’s definition snippet—a summarization box on the top of some SERPs with the text “From sources across the web”—expresses pseudo-scientific concepts. The box states that “Indigo Children are kids with indigo-colored auras”. This snippet was followed by YouTube videos promoting Indigo Children pseudoscience and a Reddit post from the /r/Psychic subreddit. Another user who may have searched only “indigo children” would have seen a snippet from Wikipedia which correctly calls indigo children pseudoscientific. However, adjacent to the snippet, the user would have seen four books for sale about indigo children. After that, the second highest-ranking return was an article from an unreliable domain identifying the “13 signs you’re an indigo child”.
@@ Methods
Due to a lack of ground truth labels for evaluating newly discovered domains, Partial F1 measures the ability of the discovery system to find unreliable domains with respect to two known lists of unreliable sources, a seed list and an evaluation list. In this case, the seed list is the PoliticalNews dataset and the evaluation list is drawn from unreliable MBFC domains (Chen and Freire 2020). As such, partial F1 measures the ability of a discovery system seeded on PoliticalNews to discover as many MBFC domains as possible (high partial recall), without discovering domains that are not in the MBFC list (high partial precision) (Castelo et al. 2019). Following Carragher, Williams, and Carley (2024a), we discuss the limitations of the partial F1 metric in more detail in Limitations and Appendix C.
@@ Results
Table 1: Mean Accuracy, F1, and standard deviations for each GNN ablation over 10 runs. $\mathcal{H}$ denotes homogeneous and $\mathcal{E}$. * denotes a social-media-only model run and evaluated on the 2,475 labeled domains mentioned in the Twitter data.
We compare our best performing model with the approach proposed in Yang and Menczer (2023) in Table 2. The authors in Yang and Menczer (2023) used ChatGPT to return the probability that 7,318 websites—7,317 of which are in our list—were unreliable labeled domains. Following the authors, we use a cutoff of 0.5 to binarize their ChatGPT predictions. Our model was trained on many of the 7,318 websites that GPT evaluated, so (full data) comparisons result in label leakage. Consequently, we provide a second (test set) evaluation, where we only compare labeled websites that our models never encountered in training. On both the full dataset and the test set evaluations, our model outperformed ChatGPT. Our model yielded an F1 that was 0.03 higher than ChatGPT on the full dataset and 0.04 higher on the test set.
@@
GPT (full data) 0.81 0.751
$\mathcal{E}_{domains+users(n2v)}$ (full data) 0.837 0.78
GPT (test set) 0.782 0.701
$\mathcal{E}_{domains+users(n2v)}$ (test set) 0.819 0.745
Table 2: Our best model outperforms the LLM approach proposed in Yang and Menczer (2023). We compare against all 7,318 websites the authors considered (full data) as well as only the websites our model never encountered during training (test set).
@@
Two annotators88825-30 year-old PhD candidates studying misinformation independently annotated the reliability of the top 20 predictions (sorted by prediction confidence) of the top performing model, heterogeneous model domains+users(n2v), over the set of all unlabeled domains. Inter-annotator agreement in the first round had a Krippendorff’s $\alpha=0.78$. The annotators then met and resolved all disagreements. We compare our model with the webgraph-based discovery (WG-BD) approach in Carragher, Williams, and Carley (2024a) [^9] As that paper’s discovery process did not rank discovered domains, we sort results by the misinformation classifier proposed in the work.. Again, two independent annotators ranked the results, yielding a Krippendorff’s $\alpha=0.69$.
@@
[…]
@@
While we attempted to mitigate limitations where possible, several substantial ones exist. First, the Twitter data were collected from 2020-2022, but the dredge words used in this paper were based on Google SERPs in 2024. We cannot say that the dredge words we extracted surfaced the same content in 2024 as when the phrases were initially used on Twitter. As with the Ukraine Biolabs example, there are also likely many cases where unreliable websites ranked highly, but were overtaken by fact checkers; our approach would not capture this. Future work could explore the impacts of better temporal alignment, which would have the added benefit of covering data voids formed around breaking news stories.
A key limitation is that the current process for extracting dredge words relies on paid third-party services. While we validated each of the extracted dredge words to ensure they are ranking in Google’s top-10, approaches to extract keyphrases from, for example, common-crawl webgraphs would provide more democratization of this line of research. We publish all dredge word data collected for this project in line with that goal. Additionally, the Partial F1 metric we use for discovery evaluation is limited, as previous work has shown many PoliticalNews and MBFC sites on which the metric relies are dead or no longer active (Carragher, Williams, and Carley 2024a). However, to our knowledge, no better metric exists for evaluating unreliable domain discovery systems. This is a needed avenue for future work.
@@ Acknowledgements
The research for this paper was supported in part by the Office of Naval Research, MURI: Persuasion, Identity, & Morality in Social-Cyber Environments under grant N000142112749 and by the Knight Foundation. It was also supported by the center for Informed Democracy and Social-cybersecurity (IDeaS) and the center for Computational Analysis of Social and Organizational Systems (CASOS) at Carnegie Mellon University. The views and conclusions are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ONR or the US Government.
@@ Appendix A Appendix A: Additional Data Details
Firstly, we drop all tweets that do not explicitly include a link to the URL. Second, users with fewer than 10 observed tweets over the time period were dropped. To further condense the dataset, we only include tweets which appeared at least 3 times in the dataset. As reposting and retweeting are important influence metrics on Twitter, this drops the tweets that likely did not receive as much attention. We also dropped domains that were only linked to by a single user and users that only linked to a single domain, as we hypothesized these pendulum nodes would be of limited use given the size of the graph. As a result of the expense of the webgraph attribute API, we chose to further restrict the Twitter data to only include tweets that mention at least one of the 43,758 domains for which we extracted attributes. This corresponds to keeping tweets that mention at least one of the 11,327 labeled domains for which we have attributes or tweets that mention one of the 11,327 and co-mention any of the 32,431 unlabeled domains.

### v3

@@ Introduction
Classifying individual texts or articles is a core problem addressed by research in misinformation detection. Such detection methods have typically relied on website content and social media data (Castelo et al. 2019; Chen and Freire 2020; Silva et al. 2021; Wang et al. 2024). Castelo et al. proposed a topic-agnostic detection system (Castelo et al. 2019) that identifies unreliable articles based on Linguistic Inquiry and Word Count features. Chen and Freire adopted the method for the task of unreliable domain discovery (Chen and Freire 2020). Their discovery system capitalized on user tendencies within the social graph, wherein a user who tweets a URL from a known unreliable source is likely to also tweet URLs from yet unknown sources. Similarly, Silva et al. combine website content and social context resulting in a misinformation page detection method that leverages heterogeneous data types (Silva et al. 2021), with a focus on early detection across a broad range of topics. The predictive power of webgraphs and social media data have been demonstrated separately in several contexts. Aswani et al. detect SEO manipulation by link-building sites by clustering on Pagerank score and Domain Authority (Aswani et al. 2021). For detecting news site bias, Aires et al. Patricia Aires, G. Nakamura, and F. Nakamura (2019) scrape cross-links from a list of prominent news sites from Media Bias Fact Check (Zandt 2022). Sehgal et al. explore a case where misinformation was spread through coordinated hyperlink and social media networks (Sehgal et al. 2021). Another recent case study demonstrated the manipulated webgraph linkages of unreliable pseudo-thinktanks (Williams and Carley 2023). Additionally, Zhang and Cabage show that social media and SEO promotion have different strengths; they find that social-sharing results in immediate but short-term boosts to traffic, while the benefits of link-building are slower to realize, but last longer (Zhang and Cabage 2017). This hints that a combination of both webgraph and social media data should be used in investigating misinformation sources. However, to our knowledge, this is the first work that attempts to combine the two contexts for unreliable domain classification or discovery.
@@
We use the domain credibility labels published by Lin et al. (2023). The authors of the work align the domain reliability ratings of all domains ranked by 6 expert groups and run imputation followed by principal component analysis to generate aggregate ratings for 11,520 domains. While the the sites are largely English-language, there are websites spanning multiple languages and target audiences, including news sites that operate in Italian, Russian, and German. We binarize the news domain rankings as ”reliable” and ”unreliable” using a threshhold of 0.5162, corresponding to the bottom two quintiles of the data. Following the methodology of Carragher, Williams, and Carley (2024a), We use the SEO toolkit service Ahrefs333ahrefs.com to extract the 10 domains which link to each of the 11,520 target domains at the highest volume (the highest-volume back-linking domains). Ahrefs feature values have been found to correlate strongly with ground-truth pagerank calculations and competing traffic estimate systems (Carragher, Williams, and Carley 2024b). In total, we extracted 43,758 domains, each with 23 domain-level attributes. Additional discussion of labels, collection details, and EDA of the webgraph can be found in Appendix A. Each edge in this dataset denotes a direct path, as a user on the source site could click a hyperlink to move to the reliability-labeled target site. We were unable to extract features for 193 domains—the domains largely appeared dead or inactive, so we excluded them from the data. This left us with 11,327 labeled domains.
@@
From this set of 3.6 billion tweets, we extract all Tweets that either link to one of our 11,327 websites or retweet, reply to, or quote a tweet containing one of our 11,327 websites. This resulted in a dataset of 320M tweets that link to over 840k unique domains. Of the original 11,327 labeled domains, only 5,504 appeared within the Twitter data. To reduce the noise in the dataset we perform several cleaning and filtration operations including dropping all tweets that do not directly link to one of our target sites. Due to length constraints, we detail Twitter cleaning in Appendix A. Following these cleaning operations, we are left with 555k users and 4.9M unique tweets which we observed tweeted or retweeted 91.2M times. This reduced dataset contains mentions of 2,475 reliability-labeled domains and 714 unlabeled backlinking domains.
@@
For the 100 websites with the respective lowest and highest reliability rankings, we extract the top-ranked $k\leq$ 1,000 Google keyphrases from Ahrefs, for a total of 34,646 keyphrases. We elected to choose a limited set of websites due to an infrastructure bottleneck: querying dredge words over 3.6B tweets is very slow on our hardware. We then used WebSearcher (Robertson and Wilson 2020) to query each of these keyphrases on Google and extracted the first 10 URLs returned in each SERP. There was no user account and cookies were not stored across queries; scraping was conducted from an IP address in Pittsburgh, Pennsylvania, which could impact search results. We kept all queries for which the target unreliable domain was returned in the top 10 Google search results. As many of the domains for which we collected dredge words did not have any queries for which they ranked in the top 10 on Google, this resulted in a set of 3,939 dredge words spanning 46 unreliable domains. We created a separate query to find mentions of these keyphrases in the 3.6B tweet covid dataset. This yielded 5.7 million tweets containing dredge words. Many of the most common mentions were explicit mentions of domain names or organization names (which can be identical to twitter handles—like “gatewaypundit”, “infowars”, and “nvic” all received over 10,000 mentions. To filter down the data, we use regular expressions to ensure each drudge word begins with a hashtag, starts a tweet, or is preceded by a white space. This retains 213 dredge words in 421k tweets that qualitatively contain less noise. Of these 421k tweets, only 9,788 (2%) explicitly linked to the unreliable domain associated with the dredge word.
@@
The user, who is encountering new information, might then be curious and compelled to search “indigo children meaning” or just “indigo children” on Google—both phrases are dredge words that surface Gaia in the top 10 results. Interestingly, after manually searching these “indigo children meaning” in June 2024, we found that Google’s definition snippet—a summarization box on the top of some SERPs with the text “From sources across the web”—expresses pseudoscientific concepts. The box states that “Indigo Children are kids with indigo-colored auras”. This snippet was followed by YouTube videos promoting Indigo Children pseudoscience and a Reddit post from the /r/Psychic subreddit. Another user who may have searched only “indigo children” would have seen a snippet from Wikipedia which correctly calls indigo children pseudoscientific. However, adjacent to the snippet, the user would have seen four books for sale about indigo children. After that, the second highest-ranking return was an article from an unreliable domain identifying the “13 signs you’re an indigo child”.
@@ Methods
Due to a lack of ground truth labels for evaluating newly discovered domains, Partial precision, recall, and F1 measure the ability of the discovery system to find unreliable domains with respect to two known lists of unreliable sources, a seed list and an evaluation list. As defined by (Chen and Freire 2020), the seed list is the PoliticalNews dataset and the evaluation list is drawn from unreliable MBFC domains:
${p=\frac{\#\text{prediction = fake {and} MBFC label = fake}}{\#\text{prediction = fake}}}$ (1)
${r=\frac{\#\text{prediction = fake {and} MBFC label = fake}}{\#\text{MBFC label = fake}}}$ (2)
${pf1=2\times\frac{p\times r}{p+r}}$ (3)
Partial F1, as given by Equation 3, measures the ability of a discovery system seeded on PoliticalNews to discover as many MBFC domains as possible (high partial recall), without discovering domains that are not in the MBFC list (high partial precision) (Castelo et al. 2019). These metrics are further limited in our DW-BD experiments, as many (53 of 74) politifact websites are either inactive or do not have keywords that rank them on Google. For those experiments, we therefore calculate partial metrics only on SERPs corresponding to dredge words from 21 active domains. We discuss the limitations of the partial F1 metric in more detail in Limitations and Appendix C.
@@ Results
Table 1: Mean Accuracy, F1, and standard deviations for each GNN ablation over 10 runs. $\mathcal{H}$ denotes homogeneous and $\mathcal{E}$ denotes heterogeneous. * denotes a social-media-only model run and evaluated on the 2,475 labeled domains mentioned in the Twitter data.
We compare our best performing model with the approach proposed in Yang and Menczer (2023) in Table 2. The authors in Yang and Menczer (2023) used ChatGPT to return the probability that 7,318 websites—7,317 of which are in our list—were unreliable labeled domains which we will denote $\mathcal{W}_{LLM}$. Following the authors, we use a cutoff of 0.5 to binarize their ChatGPT predictions. Our model was trained on many of the 7,318 websites that GPT evaluated, so comparing all data results in label leakage. Consequently, we provide a second evaluation, where we only compare labeled websites that our models never encountered in training, i.e., $\mathcal{W}_{LLM}\cap W_{test}$. On both $\mathcal{W}_{LLM}$ and the $\mathcal{W}_{LLM}\cap W_{test}$ evaluations, our model outperformed ChatGPT. Our model yielded an F1 that was 0.03 higher than ChatGPT on $\mathcal{W}_{LLM}$ and 0.04 higher on $\mathcal{W}_{LLM}\cap W_{test}$.
@@
GPT ($\mathcal{W}_{LLM}$) 0.81 0.751
$\mathcal{E}_{domains+users(n2v)}$ ($\mathcal{W}_{LLM}$) 0.837 0.780
GPT ($\mathcal{W}_{LLM}\cap\mathcal{W}_{test}$) 0.782 0.701
$\mathcal{E}_{domains+users(n2v)}$ ($\mathcal{W}_{LLM}\cap\mathcal{W}_{test}$) 0.819 0.745
Table 2: Our best model outperforms the LLM approach proposed in Yang and Menczer (2023). We evaluate against all 7,318 websites the authors considered ($\mathcal{W}_{LLM}$) and against the subset of $\mathcal{W}_{LLM}$ contained in our test set ($\mathcal{W}_{LLM}\cap\mathcal{W}_{test}$).
@@
Two annotators88825-30 year-old PhD candidates studying misinformation independently annotated the reliability of the top 20 predictions (sorted by prediction confidence) of the top performing model, heterogeneous model domains+users(n2v), over the set of all unlabeled domains. Inter-annotator agreement in the first round had a Krippendorff’s $\alpha=0.78$. The annotators then met and resolved the single disagreement. We compare our model with the webgraph-based discovery (WG-BD) approach in Carragher, Williams, and Carley (2024a) [^9] As that paper’s discovery process did not rank discovered domains, we sort results by the misinformation classifier proposed in the work.. Again, two independent annotators ranked the results, yielding a Krippendorff’s $\alpha=0.69$. Additional annotation details are discussed in Appendix D.
@@
We consider two dredge word discovery approaches: one approach (DW-BDlower and DW-BDupper) adapted from the discovery approach proposed in (Carragher, Williams, and Carley 2024a) and the other using the GNN trained with dredge word context ($\mathcal{E}_{domains+users+dredge}$), which we shorten to ($\mathcal{E}_{d+u(n2v)+dredge}$) in Table 3. The GNN without dredgewords outperforms all other baselines at every level of precision tested. We note that DW-BDlower yields higher precision than the webgraph based discovery baseline and the DW-BDupper baseline at P@5 and P@10, but the reported social media P@20 outperforms the webgraph and dredge-word discovery processes. Several of the sites returned by DW-BDupper were user-based forums, crypto exchanges, or supplement sellers, which while not necessarily reliable, are out-of-domain and not rated as unreliable.
@@
While we attempted to mitigate limitations where possible, several substantial ones exist. First, the Twitter data were collected from 2020-2022, but the dredge words used in this paper were based on Google SERPs in 2024. We cannot say that the dredge words we extracted surfaced the same content in 2024 as when the phrases were initially used on Twitter. Further analysis on the temporal alignment of keyword rankings and is warranted and would be a promising avenue for future research.
As with the Ukraine Biolabs example, there are also likely many cases where unreliable websites ranked highly, but were overtaken by fact checkers; our approach would not capture this. Future work could explore the impacts of better temporal alignment, which would have the added benefit of covering data voids formed around breaking news stories. Future work could also attempt to expand this work beyond Google.
A key limitation is that the current process for extracting dredge words relies on paid third-party services. While we validated each of the extracted dredge words to ensure they are ranking in Google’s top-10, approaches to extract keyphrases from, for example, common-crawl webgraphs would provide more democratization of this line of research. We publish all dredge word data collected for this project in line with that goal. We additionally note that due to resource and domain expert availability constraints, the discovery process was only evaluated by two domain experts; we release all annotations and annotator notes alongside the code. Additionally, the Partial F1 metric we use for discovery evaluation is limited, as previous work has shown many PoliticalNews and MBFC sites on which the metric relies are dead or no longer active (Carragher, Williams, and Carley 2024a). However, to our knowledge, no better metric exists for evaluating unreliable domain discovery systems. This is a needed avenue for future work.
@@ Acknowledgements
The research for this paper was supported in part by the Office of Naval Research, Minerva-Multi-Level Models of Covert Online Information Campaigns, Office of Naval Research under grant N000142112765, ARMY Scalable Technologies for Social Cybersecurity under grant W911NF20D0002, and by the Knight Foundation. It was also supported by the Informed Democracy and Social-cybersecurity Institute (IDeaS) and the center for Computational Analysis of Social and Organizational Systems (CASOS) at Carnegie Mellon University. The views and conclusions are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ONR or the US Government.
@@ Appendix A Appendix A: Additional Data Details
All GPU experiments experiments were run on a single NVIDIA GeForce RTX 3080. Given the small number of labels, no model took over 10 minutes to train. GNN experiments were implemented in Python3.10 using Pytorch Geometric (Fey and Lenssen 2019). We publicly release all code, annotations, and non-Twitter data.
@@
All twitter data comes from the dataset of 3.6 billion tweets described in the Data section. We first extract all tweets that link to or mention one of our 11,327 websites and all tweets that retweet, reply to, or quote a tweet containing one of those 11,327 websites. Once we have this data, we drop all users that do not explicitly include a link to one of our 11,327 target domains. Second, users with fewer than 10 observed tweets over the time period were dropped. To further condense the dataset, we only include tweets which appeared at least 3 times in the dataset—e.g., if a tweet was reposted or retweeted twice. As reposting and retweeting are important influence metrics on Twitter, this drops the tweets that likely did not receive as much attention. We also dropped domains that were only linked to by a single user and we dropped users that only linked to a single domain, as we hypothesized these pendulum nodes would be of limited use given the size of the graph. As a result of the expense of the webgraph attribute API, we chose to further restrict the Twitter data to only include tweets that mention at least one of the 43,758 domains for which we extracted attributes. This corresponds to keeping tweets that mention at least one of the 11,327 labeled domains for which we have attributes or tweets that mention one of the 11,327 and co-mention any of the 32,431 unlabeled domains.
@@ Appendix C Appendix C: Partial F1 threshhold sensitivity
Appendix D Appendix D: Annotation details
Annotators were asked to assess on a binary scale, whether each identified site was reliable or unreliable. In some of our discovered sites, this distinction is unambiguous (e.g. prominently pro-Q-anon blogs). If the website appeared to express political opinions, the annotators were instructed to search for medical or scientific claims. If authors made conspiratorial (e.g., “covid is part of a great reset”) or pseudo-scientific claims (e.g., “global warming is a hoax”), the website was judged to be unreliable. Annotations were done independently, and we report Krippendorf’s Alpha on the independently-annotated lists. In both annotation sets, annotator agreement on website reliability was substantial with $\alpha=0.78$ and $\alpha=0.69$. While we do not control for potential annotator bias, annotators never used political bias as a rationale for rating a site as unreliable. We release all annotations along with notes left by the annotators.